Operations Report – December 19th

Crew 185
Operations Report Tuesday 12/19/2017
SOL: 1
Name of person filing report: A.Passeron
Non-nominal systems:
-Generator :
The generator runned properly for the whole night.
Today, Shannon called a mechanic to definitely clean the oil circuit and fix the generator.
– Spacesuit 2 : After the engineering EVA at 12:40pm, we noticed that the ventilation systems of spacesuits N°2 and N°3 were not working. At 3:00pm, Thibault and I opened the backpack of spacesuit N°2 to repair the ventilation system. The backpack of spacesuit N°2 was full of dust. We removed the dust and the fans are now operationnal.
– Spacesuit 3 : At 3:20pm, Thibault and I opened the backpack of spacesuit N°3 to repair the ventilation system. The left fan in the backpack of spacesuit N°3 was blocked. We carefully unjammed the fan and cleaned it with blow-off canned air. It is now operationnal.
Generator (hours run):
Generator turned off by Shannon at 8:50 am.
Generator turned on by Shannon at 3:10pm.
Solar— SOC @ 7:40pm : 100%
Diesel – 80%
Propane – 58 percent volume
Ethanol Free Gasoline (5 Gallon containers for ATV) – 3.5 Gallons
Water (trailer) – 150 gallons
Water (static) – 270 gallons
Trailer to Static Pump used – No
Static to Loft Pump used – Yes
Water Meter: 50 gallons
Toilet tank emptied: Yes
ATVs Used: No
Oil Added? Yes, all the ATVs have been filled with ethanol free gasoline and are full (Except the blue Yamaha 300 without a battery)
ATV Fuel Used:
# Hours the ATVs were used today: 00:00 hours
Notes on ATVs: ATVs were nominal.
Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Opportunity rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Curiosity rover used:  No
Hours:
Beginning charge:
Ending charge:
Currently charging:
HabCar used and why, where?  No
Summary of internet: All Nominal
Summary of suits and radios: Ventilation systems of spacesuits N°2 and N°3 repaired.
Summary of Hab operations: All nominal
Summary of GreenHab operations: All nominal
Summary of ScienceDome operations: All nominal
Summary of RAMM operations: Not Operational
Summary of health and safety issues: Crew is Healthy
Questions, concerns and requests to Mission Support:

GreenHab Report – December 19th

Greenhab Report
David Murray
December 19th, 2017
Environmental control: (Choose which is appropriate and explain further if needed)
Ambient
Ambient with window/door open at 41 C from about 12:00 – 15:39
Working Hour: 15:39
Inside temp at working hour: 24 C
Outside temp during working hours: 12 C
Inside temperature high: 41 C
Inside temperature low: 22 C
Inside humidity: 16%
Heating
Functioning nominally
Turns on at 16 C and off at 20 C
Cooling
Not functional
Shade Cloth On/Off
On
The greenhab lights are on from 17:00 to 24:00 each night
Average temperature:
N/A
Changes to Crops
No changes
Morning Research Observations
N/A
Daily Water Usage for Crops
5 gallons
Time(s) of Watering Crops
15:39
Aquaponics
N/A
Narrative
The soil had thawed by this morning. Many of the tomato plants are competing for space so I began transplanting and thinning the pots. I also started Blue Curled Scotch Kale, Bloomsdale Spinach, and Golden Acre Cabbage from seed. I decided not to use the fertilizer I made yesterday and to, instead, amend my recipe and try again. I blended 1 L of hot water with 100 grams of dried moringa leaf/shoot matter, strained it through a cheese cloth, and added 5 L of water to the initial mixture. It’s in a fertilizer pump sprayer and ready to be used.
Support/Supplies Needed
N/A

Operations Report – December 19th

Crew 185
Operations Report 12/19/2017
SOL: 1
Name of person filing report: A.Passeron
Non-nominal systems:
– Heater :
The heater was not working properly on Sunday, even after reseting it. Shannon adviced us to change the air filter. On Monday morning, Thibault and I removed the air filter. It had not been changed since April 2016. At 10am, Ilaria and David went to city to buy a battery for the generator and a pack of air filters. We replaced the filter at 3:00pm. The heater is now working well.
– Generator :
On Sunday, the generator had a failure at 8:00pm, Fault message reading “Low battery”. Jump-start from Shannon’s Shuttle battery did not solve the issue. Generator had a second failure at 9:00pm on sunday, so we definitely turned it off for the night.
On Monday, Shannon helped us to change the battery at 3:00pm. We then drained oil from the radiator by pouring clean cooling liquid. The oil of the generator was also very dirty. It was so sticky that we had to start the generator for a few minutes to heat the oil and thin it. We thus were able to remove the oil filter, but only a fraction of dirty oil came out of the generator. We tried to remove as much dirty oil as possible by filling the motor with clean oil, but a valve seemed to block the oil circulation. After pouring as much clean oil as we could, we replaced the oil filter. The generator was finally turned on at 5:00pm and runned properly for the whole night.
Today, Shannon called a mechanic to definitely clean the oil circuit and fix the generator.
– Spacesuit 2 : After the engineering EVA at 12:40pm, we noticed that the ventilation systems of spacesuits N°2 and N°3 were not working. At 3:00pm, Thibault and I opened the backpack of spacesuit N°2 to repair the ventilation system. The backpack of spacesuit N°2 was full of dust. We removed the dust and the fans are now operationnal.
– Spacesuit 3 : At 3:20pm, Thibault and I opened the backpack of spacesuit N°3 to repair the ventilation system. The left fan in the backpack of spacesuit N°3 was blocked. We carefully unjammed the fan and cleaned it with blow-off canned air. It is now operationnal.
Generator (hours run):
Generator turned off by Shannon at 8:50 am.
Generator turned on by Shannon at 3:10pm.
Solar— SOC @ 7:40pm : 100%
Diesel – 80%
Propane – 58 percent volume
Ethanol Free Gasoline (5 Gallon containers for ATV) – 3.5 Gallons
Water (trailer) – 150 gallons
Water (static) – 270 gallons
Trailer to Static Pump used – No
Static to Loft Pump used – Yes
Water Meter: 50 gallons
Toilet tank emptied: Yes
ATVs Used: No
Oil Added? Yes, all the ATVs have been filled with ethanol free gasoline and are full (Except the blue Yamaha 300 without a battery)
ATV Fuel Used:
# Hours the ATVs were Used today: 00:00 hours
Notes on ATVs: ATVs were nominal.
Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Opportunity rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Curiosity rover used:  No
Hours:
Beginning charge:
Ending charge:
Currently charging:
HabCar used and why, where?  No
Summary of internet: All Nominal
Summary of suits and radios: Ventilation systems of spacesuits N°2 and N°3 repaired.
Summary of Hab operations: All nominal
Summary of GreenHab operations: All nominal
Summary of ScienceDome operations: All nominal
Summary of RAMM operations: Not Operational
Summary of health and safety issues: Crew is Healthy
Questions, concerns and requests to Mission Support:

Journalist Report – December 19th

[Sol 1]
And we have landing! This morning, we arrived on Mars as expected and we began our mission, which will last until the 30th of December with the arrival of Crew 186. But right now, we are alone in MDRS and our crew is in a great shape and a dynamic mentality.
This morning, we began the day with an engineering check to verify that the generator, the rovers, the water system… were in a good condition. It was the first time with an EVA suit for Arno and David, guided by our experienced Commander Ilaria. The station is apparently doing well and we hope that it will keep working that way until the end of the mission.
After a common lunch, we did an important meeting to set up ou schedule for the upcoming days. Then, David went to the Greehab to move his experiences forward, while our spacesuit engineers repaired two of the EVA suits that had jammed ventilators. Right now, we have six functionnal suits for five people, which is promising!
After dinner, we will do a mindfulness training and meditation and we will work on the different experiments that we will implement and run. We also hope to be able to watch a (space) movie together.
Greetings from Mars,
Thibault, ExO and spacesuit engineer for the Crew 185

Sol Summary – December 19th

Crew 185 Daily Summary Report  19/12/2017
MDRS Daily Summary Report for Sol 1
First Day of Sim
Mission Status:  Things are running smoothly.
Sol Activity Summary: We started our simulation this morning. We began our day by eating breakfast as a team and followed our meal with a mindfulness meditation session. After that, three of us put on our simulation space suits for the first time and left the hab to do an engineering check. Two crew members stayed behind for communication. When the crew returned to the hab, we ate lunch, assigned daily tasks for each crew member, and had time to work on individual projects.
Look Ahead Plan: Tomorrow, we will have our first EVA and test the X-1 space suit for the first time.
Anomalies in work:  The simulation space suit helmet fogs up easily and can drastically limit your field of vision.
David Murray crew greenhab officer and biologist
Weather: Sunny, warm
Crew Physical Status: Healthy
EVA: No EVA today
Reports to be file: EVA request
Support Requested: None